A collection of quarters and dimes is worth $6.75. The number of dimes is 4 less than three times the number of quarters. How many of each are there?
Answer by mananth(16946)

total value = 675 cents
let quarters be x
dimes = 3x-4
...
25x+10(3x-4)=675
25x+30x-40=675
55x=675+40
55x= 715
x=715/55
x= 13 quarters
dimes = 3x-4
= 3*13-4
39-4
35 dimes